a 1250 kg car is pulling a 325 trailer. together the car and trailer have an acceleration of 2.15 m/sec2. determine the net force on the car-trailer combination. what is the net force on the trailer?
1
Expert's answer
2012-07-06T07:34:58-0400
The net force would be:

F = (M+m)a = (1250+325)*2.15 = 3386.25 N
F(trailer) = ma = 325*2.15 = 698.75 N